It takes 79.91 cal to vaporize 2.0 grams of an unknown substance. What is the enthalpy of vaporization for this substance.
Answer #1

The enthalpy is calculated as J/mol or kJ/mol, or even cal/mol, kcal/mol. We have mass here, and as this is an unknown sustance, then, we cannot know the molecular weight, so we will report the enthalpy as cal/g, kcal/g and kJ/g

dH = q/n = q/m

dH = 79.91 cal / 2 = 39.955 cal/g or 0.03955 kcal/g

converting this into Joules:
39.955 * 4.184 = 167,17 J/g or 0.16717 kJ/g

Hope this helps